OpenStudy (anonymous):

Find the domain of the function. f(x) = 5x - 11

OpenStudy (anonymous):

the set of real numbers is the domain

OpenStudy (anonymous):

\[(-\infty,\infty)\]

OpenStudy (anonymous):

function is good for any value of x

OpenStudy (anonymous):

Deezil, the way you wrote it, you are including negative and positive infinity which you cannot do because we do not know what number infinity is. So it should have parenthesis around the negative and positive infinity's instead of square brackets.

OpenStudy (anonymous):

is it open or closed tyler1992?

OpenStudy (anonymous):

Open interval \[(-\infty,\infty)\]

OpenStudy (anonymous):

agreed tyler1992!

OpenStudy (anonymous):

ahh yes i mixed them up, thanks for the correction it's been a while! Of course it has to be parenthesis for \[\infty\] or inclusive value thanks for correcting that